# Localization and entanglement characterization of edge states in HgTe quantum wells in a finite strip geometry

### v2

@@ VI Conclusions
Complementary information on the structure of spin-polarization structure of near-gap states in $(k_{x},y)$ space is extracted from the RDM to the spin subsystem. Contour plots of the RDM entries show the extremal values of spin up and down and spin transfer probabilities in $(k_{x},y)$ space. Our findings indicate that the spin-polarization pattern, as shown by the diagonal elements of the RDM, is typical of the edge-state region $|k|<k_{c}$, becoming less distinct as we approach the bulk-state region $|k|>k_{c}$. Another possible criterion for differentiating edge-state from bulk-state behaviors could be this. Additionally, a study of the spin-transfer probability amplitudes (RDM off-diagonal components/coherences) concludes that, in the edge-state region $|k|<k_{c}$ (with the possible exception of a neighborhood of the $\Gamma$ point), the spin-transfer probability is significantly lower than in the bulk-state region $|k|>k_{c}$. Therefore, edge currents are more effective at preserving spin than bulk currents, one might argue.
Finally, entropies of the RDM inform on regions in $(k_{x},y)$ space where the spin sector is highly entangled with the rest of the system, due to spin-orbit coupling. Entanglement exhibits greater variability within the low-momentum $k$ region, where edge states are located, compared to the bulk state region (characterized by relatively low entropy uniformity). The behavior of the quantum correlations does not seem to depend on the particular entanglement measure used.
@@ Acknowledgments
We thank the support of Spanish MICIU through the project PID2022-138144NB-I00. OC was on sabbatical leave, from September 2023 to October 2024, at Granada University, Spain. OC thanks support from the program PASPA from DGAPA-UNAM. We would like to thank our Master’s thesis student, Alejandro Moreno, for his help at specific points in the development of the article.

### v3

@@ VII Conclusions and outlook
Complementary information on the spin-polarization structure of near-gap states in $(k_{x},y)$ space is extracted from the RDM to the spin subsystem. Contour plots of the RDM entries show the extremal values of spin up and down and spin transfer probabilities in $(k_{x},y)$ space. Our findings indicate that the spin-polarization pattern, as shown by the diagonal elements of the RDM, is typical of the edge-state region $|k|<k_{c}$, becoming less distinct as we approach the bulk-state region $|k|>k_{c}$. Another possible criterion for differentiating edge-state from bulk-state behaviors could be this. Additionally, a study of the spin-transfer probability amplitudes (RDM off-diagonal components/coherences) concludes that, in the edge-state region $|k|<k_{c}$ (with the possible exception of a neighborhood of the $\Gamma$ point), the spin-transfer probability is significantly lower than in the bulk-state region $|k|>k_{c}$. Therefore, edge currents are more effective at preserving spin than bulk currents, one might argue.
Entropies of the RDM inform on regions in $(k_{x},y)$ space where the spin sector is highly entangled with the rest of the system, due to spin-orbit coupling. Entanglement exhibits greater variability within the low-momentum $k$ region, where edge states are located, compared to the bulk state region (characterized by relatively low entropy uniformity). The behavior of the quantum correlations does not seem to depend on the particular entanglement measure used, as discussed in Section V.2 above.
Correlation functions for a real-space partition report that the edge-state region (that of higher IPR in $k$ space) has lower correlation modulus (but higher absolute correlation argument) values than the bulk state region (that of lower IPR).
If we want to compare between QI tools to distinguish the edge from the bulk states, we would say that, whereas the RDM and linear entropy entanglement measures provide better information than IPR on the spin-polarization structure of the near-gap states in $(k_{x},y)$ space, the IPR and space-correlation measures give a sharper distinction between the edge and the bulk state separating momentum region. These separating transition regions occur at different momentum cut-off values for valence and conduction bands, and they are also sensitive to the model paramenters and strip width.
Concerning future research directions, the incorporation of electron-electron interactions in the BHZ model, as already done in recent studies [43, 44], would greatly increase the physical significance of our study. We also believe that the analysis of topological phases using QI concepts and tools opens up new prospects for the development of new quantum technologies. For example, other QI concepts like “fidelity-susceptibility” turn out to capture magic twist angles arising in twisted bilayer graphene [45]. Exploring these novel quantum phenomena, termed as “twistronic” [46, 47], and others, could lead to significant opportunities for advancing quantum information devices and opening up pathways to more robust, adaptable, and efficient quantum technologies.
@@ Acknowledgments
We thank the support of Spanish MICIU through the project PID2022-138144NB-I00. OC was on sabbatical leave, from September 2023 to August 2024, at Granada University, Spain. OC thanks support from the program PASPA from DGAPA-UNAM. We would like to thank our Master’s thesis student, Alejandro Moreno, for his help at specific points in the development of the article.